# Ground Perturbation Detection via Lower-Limb Kinematic States During Locomotion

### v2

@@ Abstract
Falls during daily ambulation activities are a leading cause of injury in older adults due to delayed physiological responses to disturbances of balance. Lower-limb exoskeletons have the potential to mitigate fall incidents by detecting and reacting to perturbations before the user. Although commonly used, the standard metric for perturbation detection, whole-body angular momentum, is poorly suited for exoskeleton applications due to computational delays and additional tunings. To address this, we developed a novel ground perturbation detector using lower-limb kinematic states during locomotion. To identify perturbations, we tracked deviations in the kinematic states from their nominal steady-state trajectories. Using a data-driven approach, we further optimized our detector with an open-source ground perturbation biomechanics dataset. A pilot experimental validation with five able-bodied subjects demonstrated that our model detected ground perturbations with 97.8% accuracy and only a delay of 23.1% within the gait cycle, outperforming the benchmark by 46.8% in detection accuracy. The results of our study offer exciting promise for our detector and its potential utility to enhance the controllability of robotic assistive exoskeletons.
@@ I Introduction
The simplest approach to detect walking perturbations is to identify when the person’s body center of mass (COM) moves outside the area bounded by both feet [12, 13]. However, this detection mechanism is unsuitable for exoskeleton control due to the challenge in defining consistent boundaries throughout the gait cycle (e.g., single vs. double stance) during locomotion [14]. Whole-body angular momentum (WBAM) is commonly proposed as an alternative metric to detect perturbations during walking. [15]. WBAM represents the combined rotational momentum of all body segments around the person’s COM. Humans tightly regulate WBAM during steady-state walking; therefore, perturbations can be identified by observing a sudden increase in WBAM [16, 17].
@@
Direct tracking of lower-limb kinematic states can be a promising solution to ground perturbation detection for exoskeleton control [1]. The key anatomical points involved in walking balance, such as the foot and COM, exhibit predictable patterns during stable locomotion. Deviations in the position and velocity of these points, relative to their steady-state trajectories, serve as clear indicators of disturbances during locomotion. These deviations can be quantified with minimal time and computation using data from motion capture systems or wearable sensing [20]. Furthermore, this approach may require only 2-3 strides of gait data to capture the natural kinematic gait patterns of each new subject [21].
In this paper, we propose a novel ground perturbation detector during level-ground walkingork modeltracks 16 lower-limb kinematic states in a local coordinate system and quantifies the variation of each state from its value during steady-state walking. For our detector, we use a single threshold value to identify the perturbation onset timing across subjects and perturbation types. We evaluated our model’s detection performance in trip- and slip-type ground perturbations. These perturbation scenarios were selected because they are the most common cause of falls in aging gait during walking [11]. We hypothesize that a ground perturbation detector using lower-limb kinematic states will outperform a WBAM-based model in a faster and more accurate detection. Our underlying rationale is that the kinematic states are highly sensitive to gait disturbances, allowing efficient and accurate tracking in real-time with minimal computation [22, 23]. The development of an improved walking perturbation detector will directly support advancements in balance-assistive exoskeletons and potentially reduce the risk of falls in aging adults.
@@ II-A Open-Source Gait Perturbation Dataset
We utilized an open-source gait biomechanics dataset to investigate how individuals respond to ground translation perturbations during locomotion [24]. This dataset contained motion capture marker data, ground reaction forces, and whole-body angular momentum for 11 able-bodied subjects (4 females, age of 24.5$\pm$3.4 years, height of 175.1$\pm$7.2 cm, leg length of 91.3$\pm$5.1 cm, and body mass of 73.3$\pm$11.0 kg). The trial conditions included three sessions of 96 unique ground translation perturbations. The perturbations were varied in eight directions with three distinct magnitudes (i.e., ground translations of 5, 10, and 15 cm) using a 6 degrees of freedom actuated platform. Ground perturbations occurred at varying points within the gait cycle: 1) 50% of the double stance phase and 2) 25%, 50%, and 75% of the single stance phase. Gait phase was varied because it directly impacts subject’s response to perturbations [24]. Data from trials where subjects lost contact with the force plate during perturbation response were omitted from our analysis.
@@ II-B Tracking Kinematic States in a Local Coordinate System
To track the state of each foot, we used each subject’s heel markers. The heel position was calculated within a coordinate system centered at the global COM in 3 axes (Fig. 1A). The global COM position was calculated as the average of the positions of four pelvic markers (right/left anterior/posterior superior iliac spine markers). The relative COM position was determined by projecting the global COM vertically downward onto a coordinate system centered at the instantaneous midpoint between the feet in the horizontal plane (Fig. 1B). The vertical axis was omitted because the vertical position of the COM is minimally sensitive to ground translation perturbations [25, 12]. The velocity states were calculated by differentiating the global position over time. A dimensionality reduction technique using Principal Component Analysis (PCA) was implemented. This analysis provided a simplified visual inspection of the overall state response to perturbations (Fig. 2). It also provided qualitative validation of our prediction that the selected states would exhibit rapid and distinct deviations from steady-state behavior following a perturbation. Two different perturbations trials from the open-source dataset, one in the mediolateral direction (Fig. 2A) and one in the anteroposterior direction (Fig. 2B), were chosen for a visual evaluation.
Figure 1: Kinematic state-based perturbation detection. The relative distance between the body’s COM and each foot was calculated to track the kinematic state. A local coordinate system was established at the estimated midpoint between the feet in the horizontal plane. The COM position was then projected into this coordinate system, with a vector (red dashed lines) tracking the relative distance between the COM and the feet over time.
Figure 2: (A) Principal component analysis for a single representative trial of an (A) anteroposterior and (B) mediolateral perturbation. For both cases, perturbations with a magnitude of 5 cm occurred during the double stance phase. (C) Visual representation of $\alpha$ calculation. $\alpha$ is the euclidean distance between the current state value and the edge of 2 standard deviations from the mean using Eq. 1.
@@ II-C Quantifying Kinematic State Variance
The mean and standard deviation of each state were calculated as a function of the gait cycle using state values from 10 steady-state walking cycles. We chose 10 cycles as a conservative value for our initial development of the approach, however this number can likely be reduced to 2-3 cycles while still maintaining performance in future analysis [21]. For subsequent gait cycles, the current value of each state was subtracted from its steady-state mean value $\pm$2 standard deviations using Eq. 1,
$\alpha_{i}=\begin{aligned} &\begin{cases}0,&\text{if }|x_{i}-(\bar{x}_{i}\pm 2\sigma_{i})|=0,\\ x_{i}-(\bar{x}_{i}\pm 2\sigma_{i}),&\text{if }|x_{i}-(\bar{x}_{i}\pm 2\sigma_{i})|\neq 0.\end{cases}\end{aligned}$ (1)
where $\bar{x}$ and $\sigma$ represent the mean and standard deviation of the current state at a given time point, $\alpha$ denotes the deviation of the individual state from its natural range of variance (Fig. 2C), and $i$ indicates the current time step. During walking, $\alpha$ was assigned a value of zero if the current state value was within $\pm$2 standard deviations of the mean. A value of $\alpha$ close to zero indicated the subject’s kinematic states are within a stable basin, corresponding to steady-state walking.
Using the calculations for $\bar{x}$ and $\sigma$ above, we computed the instantaneous coefficient of variation, $C$, using Eq. 2.
$C_{i}=\frac{2\sigma_{i}}{\bar{x}_{i}}$ (2)
Using $C$ as an inverted scaling factor ensured that a state with a lower instantaneous natural variance contributed more significantly to the total stability deviation value and prevented a state with a large variance from disproportionally skewing the weighted sum [27]. Using $\alpha$ and $C$, we computed the total stability deviation value using Eq. 3,
$\varphi_{i}=\frac{1}{n}\sum_{j=1}^{n}\frac{\alpha_{i}}{C_{i}+\alpha_{i}}$ (3)
@@ II-D Threshold Optimization
To calculate the optimal threshold value for $\varphi$, which denotes a perturbation onset, we conducted a parameter sweep using the open-source perturbation dataset [24]. Values increasing by a step size of .001 from 0 to 1 were used as a threshold for perturbation detection. Detection performance was evaluated on accuracy and delay at each threshold. We determined accuracy by calculating the number of gait cycles that were correctly labeled as perturbed or unperturbed. We defined delay as the percentage of gait phase that elapsed between the onset of the perturbation and the time when $\varphi$ crossed the detection threshold. For the parameter sweep, we rewarded threshold values that minimized false detections (false positives) and undetected perturbations (false negatives), aligning with the target application of our approach for lower-limb exoskeleton control. Incorrect exoskeleton stability assistance, whether unnecessarily initiated due to false detection of a perturbation or omitted due to an undetected perturbation, carries a significant risk of user injury [28]. We rewarded reduction in detection delay second highest to ensure that fast detection was also prioritized.
@@ II-F Pilot Human Subject Experiment
Participants were exposed to a total of 20 perturbations, divided across two sessions of 10 walking trials each. In each session, five trip-type perturbations and five slip-type perturbations were randomly distributed throughout the trials. Onset timing and perturbation type were controlled and randomized using a program that remotely operated the treadmill and motion capture system (Lock Lab, Vicon, UK). This approach minimized the influence of anticipation on subjects’ responses to perturbations [24]. Each perturbation was provided no earlier than 15 seconds into each trial, ensuring that the initial mean and standard deviation of each state were calculated from at least 10 steady-state cycles. A trip perturbation was induced by an abrupt deceleration [30], during which the right treadmill belt decelerated at a rate of -3.0 m/s2 for 0.5 seconds. Conversely, a slip perturbation was induced by an abrupt acceleration in the opposite direction at 3.0 m/s2 for 0.5 seconds, such that the right belt reached -1.25 m/s. Perturbations occurred only on the right treadmill belt because gait phase was calculated based on the right heel strike. Following each perturbation, the right treadmill belt gradually (0.5 m/s2) returned to the original treadmill speed (1.25 m/s).
@@
Participants wore a set of 8 reflective markers located on left and right side at the anterior and posterior superior iliac spine, heel, and 3rd metatarsal head. Motion capture marker data and ground reaction forces were processed (Nexus 2.10.1, Vicon, UK) and analyzed using a custom Python script (Python 12.3). Marker trajectories were filtered with the Plug-In Gait Standard Woltring filter provided in the Nexus software. Trajectories with marker gaps of up to 25 frames were interpolated using a spline fill function. Using this processed data, $\varphi$ was calculated using Eqs.  (1)–(3). Heel strikes for each trial were identified using ground reaction forces from the right force plate [31]. To simulate pseudo real-time performance, a moving average was calculated for each of the 16 kinematic states over the first 10 detected gait cycles of each trial. The mean and standard deviation were continuously updated using the preceding 10 gait cycles to account for natural and gradual shifts in gait patterns. A perturbation was identified when $\varphi$ exceeded the optimized threshold. The percentage of the gait cycle elapsed between the known perturbation onset and the detection time was then calculated. Data from four trials were excluded from the analysis due to data corruption.
@@
Figure 3: (A) Experimental setup. (B) Split-belt treadmill protocol for trip- and slip-type perturbations. (C) Perturbation timing: early stance (0-10% and 90-100% of the gait cycle), mid stance (10-30% of the gait cycle), late stance (30-50% of the gait cycle), and swing phase (50-90% of the gait cycle) of the right leg. (D) Visual illustration of tracking of positional states. Blue dots represent expected positional states and red dots show actual states values during a perturbation. Black arrows indicate a distance vector between expected and actual states. Error bars indicate $\pm$1 standard deviation.
@@ III-A Optimal Perturbation Detection Threshold
A threshold for $\varphi$ of 0.125 yielded the highest model accuracy of 87.65$\pm$0.61and the smallest detection delay of 28.12$\pm$0.51of the gait cycle (Fig. 4). The upper limit of 87.65% accuracy is likely due to instances where subjects exhibited a minimal physical response to the perturbation. Our post-hoc analysis found that the kinematic states often remained within their nominal trajectories in response to 5-cm magnitude perturbations that were induced in the direction of walking. These perturbations did not significantly affect the stability of the subject. Alternatively, the upper limit of accuracy may indicate that the average of four strides of steady-state data in this dataset was insufficient to capture the natural variance [32].
@@ III-B Perturbation Detection Performance on Pilot Data
The performance of our kinematic state-based detector was evaluated on the pilot experimental dataset. Using the same optimized detection threshold, our model had a 10% improvement in accuracy on the pilot data compared to performance on the open-source dataset (Fig. 5A and 5B). In addition, perturbations were detected 5.07% of the gait cycle faster compared to the results in open-source dataset (Fig. 6).
@@ III-C Baseline Model Comparison
Implementation of the baseline WBAM-based model performed with 46.70% less accuracy than our method on the same dataset. However, it resulted in a 11.11% faster detection within the gait cycle (Fig. 6)[29]. The WBAM-based results were obtained by taking the average across sagittal and frontal planes to provide a representative analysis.

### v3

@@ Abstract
Falls during daily ambulation activities are a leading cause of injury in older adults due to delayed physiological responses to disturbances of balance. Lower-limb exoskeletons have the potential to mitigate fall incidents by detecting and reacting to perturbations before the user. Although commonly used, the standard metric for perturbation detection, whole-body angular momentum, is poorly suited for exoskeleton applications due to computational delays and additional tunings. To address this, we developed a novel ground perturbation detector using lower-limb kinematic states during locomotion. To identify perturbations, we tracked deviations in the kinematic states from their nominal steady-state trajectories. Using a data-driven approach, we further optimized our detector with an open-source ground perturbation biomechanics dataset. A pilot experimental validation with five able-bodied subjects demonstrated that our model distinguished perturbed from unperturbed gait cycles with 98.8% accuracy and only a delay of 23.1% within the gait cycle, outperforming the benchmark by 47.7% in detection accuracy. The results of our study offer exciting promise for our detector and its potential utility to enhance the controllability of robotic assistive exoskeletons.
@@ I Introduction
A simple approach to detect walking perturbations is to identify when the person’s body center of mass (COM) moves outside the area bounded by both feet [12, 13]. However, this detection mechanism is unsuitable for exoskeleton control due to the challenge in defining consistent boundaries throughout the gait cycle (e.g., single vs. double stance) during locomotion [14]. Whole-body angular momentum (WBAM) is commonly proposed as an alternative metric to detect perturbations during walking [15]. WBAM represents the combined rotational momentum of all body segments around the person’s COM. Humans tightly regulate WBAM during steady-state walking; therefore, perturbations can be identified by observing a sudden increase in WBAM [16, 17].
@@
Direct tracking of lower-limb kinematic states can be a promising solution to ground perturbation detection for exoskeleton control [1]. The key anatomical points involved in walking balance, such as the foot and COM, exhibit predictable patterns during stable locomotion. Deviations in the position and velocity of these points, relative to their steady-state trajectories, serve as clear indicators of disturbances during locomotion. These deviations can be quantified with minimal time and computation using data from motion capture systems or wearable sensing [20].
In this paper, we propose a novel ground perturbation detector during level-ground walking. Our model tracks 16 lower-limb kinematic states in a local coordinate system and quantifies the variation of each state from its value during steady-state walking. For our detector, we use a single threshold value to identify the perturbation onset timing across subjects and perturbation types. We evaluated our model’s detection performance in trip- and slip-type ground perturbations. These perturbation scenarios were selected because they are the most common cause of falls in aging gait during walking [11]. We hypothesize that a ground perturbation detector using lower-limb kinematic states will outperform a WBAM-based model by achieving faster and more accurate detection. Our underlying rationale is that the kinematic states are highly sensitive to gait disturbances, allowing efficient and accurate tracking in real-time with minimal computation [21, 22]. The development of an improved walking perturbation detector will directly support advancements in balance-assistive exoskeletons and potentially reduce the risk of falls in aging adults.
@@ II-A Open-Source Gait Perturbation Dataset
We utilized an open-source gait biomechanics dataset to investigate how individuals respond to ground translation perturbations during locomotion [23]. This dataset contained motion capture marker data, ground reaction forces, and whole-body angular momentum for 11 able-bodied subjects (4 females, age of 24.5$\pm$3.4 years, height of 175.1$\pm$7.2 cm, leg length of 91.3$\pm$5.1 cm, and body mass of 73.3$\pm$11.0 kg). The trial conditions included three sessions of 96 unique ground translation perturbations. The perturbations were varied in eight directions with three distinct magnitudes (i.e., ground translations of 5, 10, and 15 cm) using a 6 degrees of freedom actuated platform. Ground perturbations occurred at varying points within the gait cycle: 1) 50% of the double stance phase and 2) 25%, 50%, and 75% of the single stance phase. Gait phase was varied because it directly impacts subject’s response to perturbations [23]. Trials in which subjects lost force plate contact during a perturbation response, such as a jump response, were omitted from our analysis
@@ II-B Tracking Kinematic States in a Local Coordinate System
To track the state of each foot, we used the subject’s heel markers. The heel position was calculated within a coordinate system centered at the global COM in 3 axes (Fig. 1). The global COM position was calculated as the average of the positions of four pelvic markers (right/left anterior/posterior superior iliac spine markers). The relative COM position was determined by projecting the global COM vertically downward onto a coordinate system centered at the instantaneous midpoint between the feet in the horizontal plane (Fig. 1). The vertical axis was omitted because the vertical position of the COM is minimally sensitive to ground translation perturbations [24, 12]. The velocity states were calculated by differentiating the global position over time. A dimensionality reduction technique using Principal Component Analysis (PCA) was implemented. This analysis provided a simplified visual inspection of the overall state response to perturbations (Fig. 2). It also provided qualitative validation of our prediction that the selected states would exhibit rapid and distinct deviations from steady-state behavior following a perturbation. Two different perturbation trials from the open-source dataset, one in the mediolateral direction (Fig. 2A) and one in the anteroposterior direction (Fig. 2B), were chosen for a visual evaluation.
Figure 1: Kinematic state-based perturbation detection. The relative distance between the body’s COM and each foot was calculated to track the kinematic state (right). Eight relative position states and eight velocity states were tracked. A local coordinate system was established at the estimated midpoint between the feet in the horizontal plane. The COM position was then projected into this coordinate system, with a vector (red dashed lines) tracking the relative distance between the COM and the feet over time (left).
Figure 2: (A) Principal Component Analysis for a single representative trial of an (A) anteroposterior and (B) mediolateral perturbation. For both cases, perturbations with a magnitude of 5 cm occurred during the double stance phase. (C) Visual representation of state variance calculation (Eqs. 1 and 2). $\alpha$ is the Euclidean distance between the current state value and the edge of 2 standard deviations ($\sigma$) from the mean.
@@ II-C Quantifying Kinematic State Variance
The mean and standard deviation of each state were calculated as a function of the gait cycle using state values from 10 steady-state walking cycles. We chose 10 cycles as a conservative value for our initial development of the approach, however this number can likely be reduced to 2-3 cycles while still maintaining performance in future analysis [26]. For subsequent gait cycles, the current value of each state was compared to the mean and standard deviation to quantify variance using Eq. 1,
$\alpha_{i}=\begin{aligned} &\begin{cases}0,&\text{if }|\bar{x}_{i}-x_{i}|\leq 2\sigma_{i},\\ x_{i}-(\bar{x}_{i}\pm 2\sigma_{i}),&\text{otherwise}.\end{cases}\end{aligned}$ (1)
where $\bar{x}$ and $\sigma$ represent the mean and standard deviation of the current state at a given time point, $\alpha$ denotes the deviation of the individual state from its natural range of variance, and $i$ indicates the current time step. During walking, $\alpha$ was assigned a value of zero if the current state value was within $\pm$2 standard deviations of the mean. Otherwise, $\alpha$ was calculated as the euclidean distance from the current state value to 2 standard deviation above or below the mean (Fig. 2C). A value of $\alpha$ close to zero indicated the subject’s kinematic states are within a stable basin, corresponding to steady-state walking.
Next, $\alpha$ was used to calculate an instantaneous ratio of current to natural state variance (Fig. 2C). This ratio ensured that a state with a lower natural variance contributed more significantly to the total stability deviation value and prevented a state with a large natural variance from disproportionally skewing the total value. The standard deviation of each state was normalized using the coefficient of variation formula, Eq. 2a,
$C_{i}=\frac{\sigma_{i}}{\bar{x}_{i}}$ (2a)
to account for the differences in unit and scale between states [27]. $C$ represents the coefficient of variation. The ratio for each state was then computed and linearly aggregated using Eq. 2b,
$\varphi_{i}=\frac{1}{n}\sum_{j=1}^{n}\frac{\alpha_{i}}{2C_{i}+\alpha_{i}}$ (2b)
@@ II-D Threshold Optimization
To calculate the optimal threshold value for $\varphi$, which denotes a perturbation onset, we conducted a parameter sweep using the open-source perturbation dataset [23]. Values increasing by a step size of .001 from 0.5 to 4 were used as a threshold for perturbation detection. Detection performance was evaluated on accuracy and delay at each threshold. We determined accuracy by calculating the number of gait cycles that were correctly labeled as perturbed or unperturbed. We defined delay as the percentage of gait phase that elapsed between the onset of the perturbation and the time when $\varphi$ crossed the detection threshold. For the parameter sweep, we rewarded threshold values that minimized false detections (false positives) and undetected perturbations (false negatives), aligning with the target application of our approach for lower-limb exoskeleton control. Incorrect exoskeleton stability assistance, whether unnecessarily initiated due to false detection of a perturbation or omitted due to an undetected perturbation, carries a significant risk of user injury [28].
@@ II-F Pilot Human Subject Experiment
Five participants (age of 24.2$\pm$2.1 years) were exposed to a total of 20 perturbations, divided across two sessions of 10 walking trials each. Five trip-type perturbations and five slip-type perturbations were randomly distributed in each session. Participants wore a tethered safety harness to prevent injury during perturbation response (Fig. 3A). A trip perturbation was induced by an abrupt deceleration [30], during which the right treadmill belt decelerated at a rate of -3.0 m/s2 for 0.5 seconds. Conversely, a slip perturbation was induced by an abrupt acceleration in the opposite direction at 3.0 m/s2 for 0.5 seconds, such that the right belt reached -1.5 m/s (Fig. 3B). Perturbations occurred only on the right treadmill belt because gait phase was calculated based on the right heel strike. Following each perturbation, the right treadmill belt gradually (0.5 m/s2) returned to the original treadmill speed (1.25 m/s). Onset timing and perturbation type were controlled and randomized using a program that remotely operated the treadmill and motion capture system (Lock Lab, Vicon, UK). This approach minimized the influence of anticipation on subjects’ responses to perturbations and ensured even distribution of perturbation across the gait cycle (Fig. 3C). Each perturbation was induced no earlier than 15 seconds into each trial, ensuring that the initial mean and standard deviation of each state were calculated from at least 10 steady-state cycles.
@@
Participants wore a set of 8 reflective markers located on left and right side at the anterior and posterior superior iliac spine, heel, and 3rd metatarsal head. Motion capture marker data and ground reaction forces were processed (Nexus 2.10.1, Vicon, UK) and analyzed using a custom Python script (Python 12.3). Marker trajectories were filtered with the Plug-In-Gait Standard Woltring filter provided in the Nexus software. Trajectories with marker gaps of up to 25 frames were interpolated using a spline fill function. Using this processed data, $\varphi$ was calculated using Eqs. 1 and 2. Heel strikes for each trial were identified using ground reaction forces from the right force plate [31]. To simulate pseudo real-time performance, a moving average was calculated for each of the 16 kinematic states over the first 10 detected gait cycles of each trial. The mean and standard deviation were continuously updated using the preceding 10 gait cycles to account for natural and gradual shifts in gait patterns. A perturbation was identified when $\varphi$ exceeded the optimized threshold. The percentage of the gait cycle elapsed between the known perturbation onset and the detection time was then calculated. Data from four trials were excluded from the analysis due to data corruption.
@@
Figure 3: (A) Experimental setup. (B) Split-belt treadmill protocol for trip- and slip-type perturbations. (C) Perturbation timing: early stance (0-10% and 90-100% of the gait cycle), mid stance (10-30% of the gait cycle), late stance (30-50% of the gait cycle), and swing phase (50-90% of the gait cycle) of the right leg. Error bars indicate $\pm$1 standard deviation. (D) Visual illustration of tracking of positional states. Blue dots represent expected positional states and red dots show actual states values during a perturbation. Black arrows indicate a distance vector between expected and actual states.
@@ III-A Optimal Perturbation Detection Threshold
A threshold for $\varphi$ of 0.125 yielded the highest model accuracy of 87.65$\pm$0.61and the smallest detection delay of 28.12$\pm$0.51of the gait cycle (Fig. 4). The upper limit of 87.65% accuracy is likely due to instances where subjects exhibited a minimal physical response to the perturbation. Our post-hoc analysis found that the kinematic states often remained within their nominal trajectories in response to 5-cm magnitude perturbations that were induced in the direction of walking.
@@ III-B Perturbation Detection Performance on Pilot Data
The performance of our kinematic state-based detector was evaluated on the pilot experimental dataset using the same optimized detection threshold (Fig. 5A). The model classified gait cycles in the pilot dataset as either perturbed or unperturbed with 98.8% accuracy (Fig. 5B). In addition, perturbations were detected 5.07% of the gait cycle faster compared to the results in open-source dataset (Fig. 6).
@@ III-C Baseline Model Comparison
Implementation of the baseline WBAM-based model performed with 41.0% less accuracy than our method for the open-source dataset. However, it resulted in a 11.11% faster detection within the gait cycle. The WBAM-based results were obtained by taking the average across sagittal and frontal planes to provide a representative analysis (Fig. 6) [29].